OpenStudy (anonymous):

A cannery processed 605 pounds of strawberries in 3.5 hours. The cannnery processed 2100 pounds in 10 hours. Write a linear equation to model the weight of strawberries "S" processed in "T" hours. How many pounds of strawberries can be processed in 12 hours?

OpenStudy (anonymous):

2560 strawberries in 12 hours(:

OpenStudy (saifoo.khan):

We basically have two points. (3.5, 605) and (10,2100) Solve them to find the equation of the line.

OpenStudy (anonymous):

\[\frac{1495}{6.5}(x-3.5)=y-605\]

OpenStudy (saifoo.khan):

1495/6.5 thing took me a minute to understand. Hehe.

OpenStudy (anonymous):

lol yeah thats what i got, so it would just be y=1495/6.5x+b?

OpenStudy (saifoo.khan):

y = 230x + b Would be better. And S = 230T -200 will mark your answer as correct.

OpenStudy (anonymous):

So is the answer 2254, 2560, 2960, or 1874?